OpenStudy (anonymous):

Indefinite integral wrt x of 1 over (1+x^2)^2 (Take x = Tan u)

OpenStudy (anonymous):

\[\int\limits_{}^{}1/(1+x ^{2})^{2}\]

OpenStudy (anonymous):

Shouldn't it be ∫1/(1+x2)2dx ?

OpenStudy (amistre64):

1+tanu ^2 = sec^2 tanu=x sec^2 = dx

OpenStudy (anonymous):

\[If x = \tan x, \int\limits\limits 1/(1+\tan^2x)^2dx = \int\limits\limits 1/\sec^2xdx = \int\limits\limits \cos^2 xdx = -\sin^2x +C \] Did I miss anything?

OpenStudy (amistre64):

sec^2/sec^4 = 1/sec^2 = cos^2

OpenStudy (anonymous):

the derivative of tanx

OpenStudy (anonymous):

-sin^2x+C

OpenStudy (amistre64):

cant tell; it ran off the screen

OpenStudy (amistre64):

there should be a tan-1(x) invloved in the integral I believe

OpenStudy (anonymous):

Here's the whole expression: If x=tanx,∫1/(1+tan2x)2dx=∫1/sec2xdx=∫cos2xdx=−sin2x+C

OpenStudy (anonymous):

That's not a 2*x in the expression. Should be tan^2 x

OpenStudy (anonymous):

Tangent you made a mistake early on and carried it through all your work. You didn't substitute dx=sec^2 u du

OpenStudy (anonymous):

Thanks for that link, Amistre. And thank you for pointing that out, Chaguanas. Can you illustrate the correct, step-by-step solution? (This is a fun problem.)

OpenStudy (amistre64):

cos^2(u) = 1/2 + cos(2u)/2 ; integrate that

OpenStudy (anonymous):

I wonder: Why do you need to substitute not only for x, but also for dx?

OpenStudy (amistre64):

when you change variables; you have to calibrate for it

OpenStudy (amistre64):

you cant just up and alter equation; yo have to make them equal

OpenStudy (anonymous):

Tangent, it is the process called u-substitution. You are thinking about algebraic substitution which is something else.

OpenStudy (amistre64):

the dx part carries information that needs to be adjust for

OpenStudy (amistre64):

substitutions are like packing up the house and moving down the street; but if you dont modify your dx part, its like leaving the dog behind :)

OpenStudy (anonymous):

(This is SO much better than studying alone!!!) I noticed estudier's clue that x should be set = to tan x, and I tried to employ trigonometric substitution. Admittedly I haven't had much practice in these types of integrals, so thanks for helping me along. Question: If you set x = tan^2 u, dx = sec^2 u du, then doesn't this happen?\[\int\limits 1/(1+\tan^2u)^2*\sec^2udu=\int\limits1/(\sec^2u)^2*\sec^2udu=\int\limits1/\sec^2u*du?\] P.S. Good analogy, Amistre.

OpenStudy (anonymous):

(Gah!) \[\int\limits1/\sec^2u du\] This is the last term.

OpenStudy (anonymous):

Yes. That is what you integrate.

OpenStudy (anonymous):

I think Amistre64 png is mostly right, bit off towards the end.

OpenStudy (anonymous):

Using the previous substitution dx=sec^2 u du, du = dx/sec^2u. ∫1/sec2udu = ∫cos^2udu = ∫cos^2u/sec^2u dx = ∫cos^4udx But maybe you don't need to go that far, integrating with respect to x with a function that involves u?

OpenStudy (anonymous):

Where I'm mostly stuck is getting to arctan. That part doesn't make too much sense (maybe because I didn't pay enough attention to integration tables?)

OpenStudy (amistre64):

wolfram did it :) I felt lazy and didnt want to retype all of it :)

OpenStudy (anonymous):

Tangent just leave it at\[\int\limits_{}^{}\cos ^{2}u\]Convert this with a double angle formula which is easy to integrate.

OpenStudy (anonymous):

@ amistre64 I am only going to post integrals that Wolfram can't do from now on:-)

OpenStudy (anonymous):

Oh no, you don't! (D:) In all honesty Wolfram really helped. Now to brush up on my trig

OpenStudy (anonymous):

Hmm..nobody seems too keen on this one http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e049a8f0b8b35337e2a1700